### v1

# Coincidence detection techniques for direct measurement of many-body correlations in strongly correlated electron systems

Abstract
Research on strongly correlated electron systems faces a fundamental challenge due to the complex nature of intrinsic many-body correlations. A key strategy to address this challenge lies in advancing experimental methods that can directly probe and elucidate the underlying many-body correlations. In this perspective article, we discuss the theoretically proposed coincidence detection techniques, which are designed to directly measure two-body correlations in various particle-particle and particle-hole channels, with momentum, energy, and/or spatial resolution. We also explore the prospects of these coincidence detection techniques for future theoretical and experimental developments. The successful implementation and refinement of these coincidence detection techniques promise to deliver powerful new approaches for unraveling long-standing puzzles in strongly correlated electron systems, such as the enigmatic mechanism of unconventional superconductivity and the long-sought quantum spin liquids. Furthermore, these coincidence detection techniques will offer powerful new methods to investigate novel phenomena like itinerant magnetism and electronic nematicity in quantum materials.
I Introduction
Strongly correlated electron systems represent a major focus of modern condensed matter physics due to their novel physical properties. Systems such as unconventional superconductors (Anderson, 1987; Lee et al., 2006; Chen et al., 2014; Stewart, 2011, 1984, 2001) and quantum magnets (Balents, 2010; Zhou et al., 2017; Chamorro et al., 2021) exhibit a range of unconventional and anomalous phenomena, including non-BCS superconductivity, pseudogap physics, strange metal behavior, electronic nematicity, quantum spin liquids, and quantum phase transitions. These emergent properties are governed by intrinsic many-body correlations that can not be adequately described within well established theoretical frameworks like the perturbative Fermi liquid theory or the Ginzburg–Landau-Wilson theory of phase transitions. The development of new theoretical and experimental methods to study these correlation-driven phenomena remains a significant challenge. In this perspective article, we focus on recent progress and ongoing efforts to develop new experimental techniques within this research field.
A variety of experimental techniques have been developed to investigate the superconducting states of unconventional superconductors. Zero electric resistance and the diamagnetic Meissner effect are two key signatures of the superconducting phase, while specific heat measurements establish its macroscopic phase transition nature. Evidence for Cooper-pair-mediated superconductivity comes from flux quantization experiments (Gough et al., 1987). The superfluid density is accessible via measurements of the magnetic penetration depth (Panagopoulos and Xiang, 1998), and the low-energy optical conductivity reveals the electrodynamic response of the superfluid condensate and the corresponding charge-carrier excitations (Basov and Timusk, 2005). For pairing symmetry, angle-resolved photoemission spectroscopy (ARPES) (Damascelli et al., 2003) and scanning tunneling spectroscopy (STS) (Fischer et al., 2007) offer momentum and spatial resolutions, respectively. Further insights are provided by nuclear magnetic resonance (NMR) (Asayama et al., 1991) and electronic Raman spectroscopy (Devereaux and Hackl, 2007). Phase-sensitive techniques have been especially critical in identifying the pairing symmetry of high-Tc cuprate superconductors (Tsuei and Kirtley, 2000). Together, these methods enable a multifaceted investigation of superconductivity across different channels and scales. However, a fundamental limitation remains that none directly probes the intrinsic two-body correlations of Cooper pairs. Consequently, they can not unequivocally determine the microscopic pairing mechanism responsible for high-Tc superconductivity.
Beyond the mystery of high-Tc superconductivity itself, the parent normal states in unconventional superconductors exhibit equally esoteric properties. In cuprate superconductors, a pseudogap phase manifests in transport, thermodynamic, and magnetic channels (Li, 2021; Luo et al., 2008; Proust and Taillefer, 2019), while the strange metal state is characterized by an enigmatic linear-$T$ resistivity (Keimer et al., 2015; Li, 2021; Proust and Taillefer, 2019; Varma, 2020; Zaanen, 2019). Similarly, electronic nematicity is ubiquitous in iron-based superconductors, with ARPES measurements revealing a d-wave symmetric feature within the nematic phase (Yi et al., 2011; Su et al., 2015). Heavy-fermion systems, in turn, display non-Fermi liquid behaviors and diverse quantum critical phenomena, as observed in specific heat, magnetic susceptibility, and resistivity measurements (Stewart, 2001). These phenomena collectively point to strong many-body correlations that remain elusive to direct experimental probes.
A particularly intriguing class of quantum states in strongly correlated electron systems is the theoretically proposed quantum spin liquids (Anderson, 1987). The investigation of these states remains a prominent and active frontier in condensed matter physics (Balents, 2010; Zhou et al., 2017; Savary and Balents, 2016; Chamorro et al., 2021). Experimentally, techniques such as muon spin spectroscopy, inelastic neutron scattering (INS), NMR, and thermal transport measurements are employed to study their magnetic, thermodynamic, and transport properties (Savary and Balents, 2016). Theoretically, models based on spin fractionalization predict the emergence of fractionalized excitations in these systems (Lee et al., 2006; Auerbach, 1994; Kitaev, 2006; Gao and Chen, 2020). However, such fractionalized excitations do not directly couple to conventional experimental probes, thereby presenting a major challenge for their direct detection and characterization.
As outlined in the above brief review, a wide array of experimental techniques has been employed to investigate strongly correlated electron systems. Despite the substantial body of experimental data thus accumulated, research in this field appears to be approaching a bottleneck. This impasse may be attributed to two principal factors. First, the essential intrinsic signatures have yet to be systematically distilled from the vast and complex experimental data, hindering the development of sound phenomenological theories. Second, the existing experimental approaches may be fundamentally limited in their ability to probe the intrinsic many-body correlations and their physical manifestations. Unraveling the foundational correlation-driven physics will likely require the development of new experimental theories and advanced technologies.
The intrinsic $N$-body correlations in strongly correlated electron systems can be accessed via $N$-body response measurements. For example, the intrinsic two-body correlations can be directly probed via two-body response functions which are dominated by the second-order perturbation of the interaction between the target material and an external probe field. Guided by this principle, several coincidence detection techniques have recently been proposed for directly measuring two-body correlations in strongly correlated electron systems. Notable examples include coincidence angle-resolved photoemission spectroscopy (cARPES), which is designed to access two-body correlations in the particle-particle channel, thereby offering a new pathway to elucidate the mechanism of unconventional superconductivity (Su and Zhang, 2020; Devereaux et al., 2023; Cao and Su, 2025). Meanwhile, coincidence inelastic neutron scattering (cINS) has been proposed to probe two-spin correlations, opening avenues for studying exotic quantum spin liquids (Su et al., 2021; Cao and Su, 2025). More recently, a coincidence double-tip scanning tunneling spectroscopy (STS) method has also been introduced, enabling spatially resolved detection of dynamical two-body correlations (Su et al., 2025). Other proposed coincidence detection techniques include coincidence angle-resolved photoemission and inverse-photoemission spectroscopy (cARP/IPES) in the particle-hole channel for studying itinerant magnetism and electronic nematicity (Su and Zhang, 2020), coincidence angle-resolved inverse-photoemission spectroscopy (cARIPES) in the particle-particle channel for correlated electrons above the Fermi energy (Su and Zhang, 2020), and double photoemission (DPE) for exploring center-of-mass physics of two-body correlations (Berakdar, 1998; Su et al., 2024). The fundamental principles underlying these coincidence detection techniques can be summarized as follows (Su et al., 2024): (1) Intrinsic two-body correlations are directly captured through two-body coincidence detection, and (2) the coincidence detection probability is governed by the second-order perturbation of the interaction between the target material and the external probe field.
In this perspective article, we will explore the conceptual foundations of these coincidence detection techniques, covering their key theoretical formulations and applications for probing two-body correlations in various specific channels. We will also address the challenges and prospects for their theoretical development and experimental implementation.
II Coincidence Detection: Fundamental Principles
We begin with the Hamiltonian of the combined system comprising the target electron system and the external probe field,
$H=H_{0}+V,$ (1)
where $H_{0}$ includes the individual Hamiltonians of the two subsystems, and $V$ describes the electron-probe interaction terms involving single-electron scattering processes. Let us introduce the time-evolution $S$-matrix defined as
$S(t,t_{i})=T_{t}\exp\left[-\frac{i}{\hbar}\int_{t_{i}}^{t}dt_{1}V_{I}(t_{1})\right],$ (2)
where $T_{t}$ is a time ordering operator, and $V_{I}$ denotes the representation of $V$ in the interaction picture:
$V_{I}(t)=e^{iH_{0}t/\hbar}Ve^{-iH_{0}t/\hbar}.$ (3)
Here, $H_{0}$ is assumed to be time-independent.
Let the initial states of the combined system at time $t_{i}$ be described by $|\Phi_{I}\rangle$} with statistical probabilities $P_{I}$, where $I$ labels the state index. The first-order perturbation of the $S$-matrix, corresponding to single-electron scattering processes, is given by
$S^{(1)}(t,t_{i})=-\frac{i}{\hbar}\int_{t_{i}}^{t}dt_{1}V_{I}(t_{1}).$ (4)
The associated single-electron scattering probability reads
$\Gamma^{(1)}=\sum_{IF}P_{I}\left|\langle\Phi_{F}|S^{(1)}(t_{f},t_{i})|\Phi_{I}\rangle\right|^{2},$ (5)
where $|\Phi_{F}\rangle$} are the final states at time $t_{f}$, labeled by index $F$. Similarly, the second-order perturbation expansion takes the form
$S^{(2)}(t,t_{i})=\frac{1}{2}\left(-\frac{i}{\hbar}\right)^{2}\iint_{t_{i}}^{t}dt_{2}dt_{1}T_{t}[V_{I}(t_{2})V_{I}(t_{1})],$ (6)
which captures two-electron scattering processes via the time-ordered product $T_{t}[V_{I}(t_{2})V_{I}(t_{1})]$. The corresponding coincidence detection probability of two-electron scattering processes is defined by
$\Gamma^{(2)}=\sum_{IF}P_{I}\left|\langle\Phi_{F}|S^{(2)}(t_{f},t_{i})|\Phi_{I}\rangle\right|^{2}.$ (7)
While $\Gamma^{(1)}$ involves single-electron scattering processes and thus probes the dynamical single-electron physics of the target electron system, $\Gamma^{(2)}$ arises from two-electron scattering processes and captures the dynamical two-body correlations. This constitutes the conceptual foundation of coincidence detection, which uses the measurement of two-electron scattering processes to reveal the intrinsic two-body correlations in the target electron system. The fundamental principles underlying coincidence detection can be formally stated as follows (Su et al., 2024):
  1. 1.
    Intrinsic two-body correlations are directly captured through two-body coincidence detection.
  2. 2.
    The coincidence detection probability is governed by the second-order perturbation of the interaction between the target material and the external probe field.
These principles provide a unified framework for the variety of coincidence detection techniques introduced in recent works (Su and Zhang, 2020; Su et al., 2021; Devereaux et al., 2023; Cao and Su, 2025; Su et al., 2025).
Beyond the coincidence detection probability $\Gamma^{(2)}$, coincidence detection can also be formulated in terms of operator expectations. Consider an operator $O$ describing single-electron scattering processes. Its expectation value is given by
$\overline{\langle O(t)\rangle}=\langle S(t_{i},t)O_{I}(t)S(t,t_{i})\rangle,$ (8)
where $\langle A\rangle\equiv\operatorname{Tr}[\rho_{0}A]$ with $\rho_{0}=\frac{1}{Z}\exp(-\beta H_{0})$, and $S(t_{i},t)=[S(t,t_{i})]^{\dagger}$. Introducing the contour-time ordered $S$-matrix,
$S_{c}(t_{i},t_{i})=T_{c}[S(t_{i},t)S(t,t_{i})],$ (9)
where $T_{c}$ is the contour-time ordering operator along the contour $t_{i}\rightarrow t\rightarrow t_{i}$, we may write $S_{c}$ explicitly as
$S_{c}(t_{i},t_{i})=T_{c}\exp\left[-\frac{i}{\hbar}\int_{c}dt_{1}V_{I}(t_{1})\right],$ (10)
with the integration performed along the closed time contour. The expectation value of $O$ then becomes
$\overline{\langle O(t)\rangle}=\langle T_{c}S_{c}(t_{i},t_{i})O_{I}(t)\rangle.$ (11)
For two operators $O_{1}$ and $O_{2}$, each describing single-electron scattering processes, the coincidence detection expectation value is defined by
$\overline{\langle O_{1}(t)O_{2}(t)\rangle}=\langle T_{c}S_{c}(t_{i},t_{i})O_{1I}(t)O_{2I}(t)\rangle.$ (12)
This expression provides the fundamental formulation for coincidence detection in the context of coincidence double-tip STS measurements (Su et al., 2025).
III Coincidence Detection Techniques
We will explore the coincidence detection techniques in this section, which are proposed based upon the fundamental principles summarized in Sec. II.
III.1 cARPES
Coincidence detection of two photoelectric processes, in which two incident photons emit two photoelectrons from the target material, captures two-body correlations of the correlated electrons in the particle-particle channel. — cARPES
In cARPES coincidence detection (schematically illustrated in Fig. 1 (a) of Reference (Cao and Su, 2025)), two incident photons induce two photoelectric processes. Two single-photoelectron detectors separately detect the emitted photoelectrons. A coincidence detector records the simultaneous occurrence events of two photoelectric processes. The corresponding coincidence detection probability carries information on two-body correlations of the target electron system.
Suppose the two photoelectric processes are governed by two electron-photon interaction terms
$V_{1}$ $=$ $g_{1}d^{{\dagger}}_{\mathbf{k}_{1}+\mathbf{q}_{1}\sigma_{1}}c_{\mathbf{k}_{1}\sigma_{1}}a_{\mathbf{q}_{1}\lambda_{1}},$
$V_{2}$ $=$ $g_{2}d^{{\dagger}}_{\mathbf{k}_{2}+\mathbf{q}_{2}\sigma_{2}}c_{\mathbf{k}_{2}\sigma_{2}}a_{\mathbf{q}_{2}\lambda_{2}},$ (13)
where $a_{\mathbf{q}\lambda}$ annihilates an incident photon with momentum $\mathbf{q}$ and polarization $\lambda$, $c_{\mathbf{k}\sigma}$ annihilates a target electron with momentum $\mathbf{k}$ and spin $\sigma$, and $d^{{\dagger}}_{\mathbf{k}\sigma}$ creates a photoelectron. The coupling constants $g_{1}\equiv g(\mathbf{k}_{1};\mathbf{q}_{1}\lambda_{1})$ and $g_{2}\equiv g(\mathbf{k}_{2};\mathbf{q}_{2}\lambda_{2})$ characterize the interaction strengths.
The coincidence detection probability for the simultaneous occurrence of two photoelectric processes is given by (Su and Zhang, 2020)
$\Gamma^{(2)}=\frac{1}{Z}\sum_{\alpha\beta}e^{-\beta E_{\alpha}}\left\lvert\langle\Phi_{\beta}\rvert S^{(2)}(t_{f},t_{i})\rvert\Phi_{\alpha}\rangle\right\rvert^{2},$ (14)
with initial states $\lvert\Phi_{\alpha}\rangle=\lvert\Psi_{\alpha}\rangle\otimes\lvert 1_{\mathbf{q}_{1}\lambda_{1}}1_{\mathbf{q}_{2}\lambda_{2}}\rangle_{p}\otimes\lvert 0\rangle_{d}$ and final states $\lvert\Phi_{\beta}\rangle=\lvert\Psi_{\beta}\rangle\otimes\lvert 0\rangle_{p}\otimes\lvert 1_{\mathbf{k}_{1}+\mathbf{q}_{1}\sigma_{1}}1_{\mathbf{k}_{2}+\mathbf{q}_{2}\sigma_{2}}\rangle_{d}$. Here $\lvert\Psi_{\alpha}\rangle$ and $\lvert\Psi_{\beta}\rangle$ are eigenstates of the target electron system with energies $E_{\alpha}$ and $E_{\beta}$, respectively. The relevant second-order $S$-matrix term reads
$S^{(2)}(t_{f},t_{i})=\left(-\frac{i}{\hbar}\right)^{2}\iint^{t_{f}}_{t_{i}}dt_{2}dt_{1}T_{t}\left[V_{2I}(t_{2})V_{1I}(t_{1})\right].$ (15)
Let us introduce a two-body Bethe-Salpeter wave function in the particle-particle channel as (Gell-Mann and Low, 1951; Salpeter and Bethe, 1951)
[…]
In the limit $t_{i}\to$-$\infty$ and $t_{f}\to+\infty$, the coincidence detection probability becomes (Su and Zhang, 2020)
$\Gamma^{(2)}=\frac{\left(g_{1}g_{2}\right)^{2}}{\hbar^{4}}\frac{1}{Z}\sum_{\alpha\beta}e^{-\beta E_{\alpha}}\big|\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1},\mathbf{k}_{2}\sigma_{2};\Omega_{c},\omega_{r}\right)\big|^{2}.$ (17)
Here we have identified $\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1},\mathbf{k}_{2}\sigma_{2};t_{c},t_{r}\right)=\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1}t_{1};\mathbf{k}_{2}\sigma_{2}t_{2}\right)$, where $t_{c}=\frac{1}{2}(t_{1}+t_{2})$ and $t_{r}=t_{2}-t_{1}$ denote the center-of-mass and relative times, respectively. The corresponding Fourier transformations are defined by
$\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1},\mathbf{k}_{2}\sigma_{2};t_{c},t_{r}\right)$
$=$ $\frac{1}{\left(2\pi\right)^{2}}\iint_{-\infty}^{+\infty}d\Omega d\omega\,\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1},\mathbf{k}_{2}\sigma_{2};\Omega,\omega\right)e^{-i\Omega t_{c}-i\omega t_{r}},$
$\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1},\mathbf{k}_{2}\sigma_{2};\Omega,\omega\right)$
$=$ $\iint_{-\infty}^{+\infty}dt_{c}dt_{r}\,\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1},\mathbf{k}_{2}\sigma_{2};t_{c},t_{r}\right)e^{i\Omega t_{c}+i\omega t_{r}}.$
The center-of-mass frequency $\Omega_{c}$ and the inner-pair relative frequency $\omega_{r}$ in $\Gamma^{(2)}$ are defined as
$\Omega_{c}=\frac{1}{\hbar}(E_{1}^{(2)}+E_{2}^{(2)}),\ \omega_{r}=\frac{1}{2\hbar}(E_{2}^{(2)}-E_{1}^{(2)}),$ (18)
with $E_{1}^{(2)}$ and $E_{2}^{(2)}$ being the transferred energies in the two photoelectric processes
$E^{(2)}_{1}=\varepsilon^{(d)}_{\mathbf{k}_{1}+\mathbf{q}_{1}\sigma_{1}}+\Phi-\hbar\omega_{\mathbf{q}_{1}},$
$E^{(2)}_{2}=\varepsilon^{(d)}_{\mathbf{k}_{2}+\mathbf{q}_{2}\sigma_{2}}+\Phi-\hbar\omega_{\mathbf{q}_{2}}.$ (19)
Here $\varepsilon^{(d)}$ are the photoelectron energies, $\hbar\omega$ are the photon energies, and $\Phi$ is the target-material work function.
It is noted that the Bethe-Salpeter wave function $\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1}t_{1};\mathbf{k}_{2}\sigma_{2}t_{2}\right)$ describes the dynamical annihilation of two electrons of the target material. The coincidence detection probability $\Gamma^{(2)}$, which probes the frequency-space wave function $\phi^{(2)}_{\alpha\beta}(\mathbf{k}_{1}\sigma_{1},\mathbf{k}_{2}\sigma_{2};\Omega_{c},\omega_{r})$, therefore directly reveals dynamical two-body correlations in the target electron system. Furthermore, this approach enables explicit resolution of both center-of-mass and inner-pair relative frequency dependencies inherent to these dynamical two-body correlations (Su and Zhang, 2020).
In the pairing channel for Cooper pairs, the inner-pair relative dynamics is governed by a time-retarded attractive interaction, such as the phonon-mediated pairing interaction in conventional metallic superconductors. Therefore, the inner-pair relative dynamics encoded in the Bethe-Salpeter wave function obtained from cARPES can shed light on the microscopic pairing interaction governing the formation of Cooper pairs. This makes cARPES a vital tool for uncovering the microscopic pairing mechanism of superconductivity (Su and Zhang, 2020; Devereaux et al., 2023; Cao and Su, 2025; Huang et al., 2021; Sobota et al., 2021; Alexandradinata et al., 2025). In cuprate superconductors, the microscopic origin of superconductivity has remained a long-standing puzzle. Does it arise from the coherent motion of holons with paired spinons, as suggested by the resonating valence bond (RVB) theory (Anderson, 1987; Lee et al., 2006)? Or is it mediated by a pairing “glue” such as dynamical spin fluctuations in a nearly antiferromagnetic Fermi liquid (Monthoux and Pines, 1993, 1994)? Could it even be related to a Feshbach resonance scenario (Bohrdt et al., 2023; Homeier et al., 2024, 2025)? Measurements via cARPES in the pairing channel hold the potential to resolve this fundamental question. Moreover, cARPES is poised to make significant contributions to identifying the microscopic pairing mechanisms in other classes of unconventional superconductors, including iron-based superconductors (Chen et al., 2014; Stewart, 2011) and heavy-fermion superconductors (Stewart, 1984, 2001). It has also been proposed recently as a promising technique for probing the fluctuating nature of light-enhanced d-wave superconductivity (Wang et al., 2021), as well as the off-diagonal matrix-element driven spin spectra in topological states (Lin et al., 2024).
In the initial experimental design for cARPES measurement (Su and Zhang, 2020; Cao and Su, 2025), a coincidence detector is employed to instantaneously record the simultaneous occurrence of two photoelectric processes. This configuration is termed instantaneous cARPES. More recently, an alternative implementation was proposed (schematically shown in Fig. 1 (b) of Reference (Cao and Su, 2025)), termed post-experiment cARPES (Cao and Su, 2025). To capture the dynamical two-body correlations in the target electron system, the time interval between the initiations of two photoelectric processes must be shorter than the characteristic time scale $t_{c}$ of the physics under investigation. This can be achieved using a pulse photon source, where both incident photons originate from the same pulse with a time width $\Delta t_{p}\leq t_{c}$. Furthermore, to clearly distinguish individual coincidence detection events, the time interval $\Delta t_{d}$ between consecutive photon pulses must satisfy $\Delta t_{d}\gg\Delta t_{p},t_{c}$. A further refinement in the post-experiment cARPES is the use of two counting recorders $R_{1}$ and $R_{2}$, which record photoelectrons arriving at two detectors $D_{1}$ and $D_{2}$, respectively. For the $n$-th photon pulse, the recorded photoelectron numbers $a_{1}(n)$ and $a_{2}(n)$ in $R_{1}$ and $R_{2}$ take values of 0 or 1. The coincidence counting for the $n$-th pulse is then computed as $a_{1}(n)\times a_{2}(n)$ during post-processing. This approach, without the coincidence detector used in the instantaneous cARPES, enables easier and more efficient identification of two-photoelectric-process coincidence events. As a result, the post-experiment cARPES technique offers a highly efficient and powerful means for measuring two-body correlations in the target electron system (Cao and Su, 2025).
It is worth noting that the cARPES technique shares conceptual similarities with time-resolved ARPES measurements of two-electron emission events, which was proposed to investigate electronic noise correlations in materials (Stahl and Eckstein, 2019). The related theoretical framework has also been developed to probe two-electron interactions via coincidence detection of two photoelectric processes (Kemper et al., 2025).
III.2 cINS
Coincidence detection of two neutron-scattering processes is essentially a measurement of two-spin correlations in the target magnetic material. — cINS
A coincidence detection technique in the spin-spin channel, termed cINS, can be analogously proposed (Su et al., 2021). Here, the two incident particles are neutrons, which interact with the local spin moments of the target magnetic material and become scattered. These scattered neutrons can be coincidentally detected either by a dedicated coincidence detector (Su et al., 2021) or via a post-experiment coincidence counting method (Cao and Su, 2025).
The two neutron-scattering processes are governed by the following electron-neutron interaction terms (Lovesey, 1984; Squries, 1996; Fernandez-Alonso and Price, 2013; Su et al., 2021; Cao and Su, 2025)
$V_{1}$ $=$ $\sum_{\sigma_{i_{1}}\sigma_{f_{1}}}g(\mathbf{q}_{1})f^{{\dagger}}_{\mathbf{q}_{f_{1}}\sigma_{f_{1}}}\boldsymbol{\tau}_{\sigma_{f_{1}}\sigma_{i_{1}}}f_{\mathbf{q}_{i_{1}}\sigma_{i_{1}}}\cdot\mathbf{S}_{\perp}(\mathbf{q}_{1}),$
$V_{2}$ $=$ $\sum_{\sigma_{i_{2}}\sigma_{f_{2}}}g(\mathbf{q}_{2})f^{{\dagger}}_{\mathbf{q}_{f_{2}}\sigma_{f_{2}}}\boldsymbol{\tau}_{\sigma_{f_{2}}\sigma_{i_{2}}}f_{\mathbf{q}_{i_{2}}\sigma_{i_{2}}}\cdot\mathbf{S}_{\perp}(\mathbf{q}_{2}).$ (20)
[…]
Following a procedure analogous to that of cARPES, we can introduce the second-order term of the time-evolution $S$-matrix corresponding to the two electron-neutron interaction terms, enabling the calculation of the coincidence detection probability $\Gamma^{(2)}$ for the simultaneous occurrence of two neutron-scattering processes. Let us define a two-spin Bethe-Salpeter wave function as (Su et al., 2021; Cao and Su, 2025)
$\phi^{(ij)}_{\alpha\beta}(\mathbf{q}_{1}t_{1},\mathbf{q}_{2}t_{2})=\langle\Psi_{\beta}|T_{t}S^{(j)}_{\perp}(\mathbf{q}_{2},t_{2})S^{(i)}_{\perp}(\mathbf{q}_{1},t_{1})|\Psi_{\alpha}\rangle,$ (21)
where $i$ and $j$ label the Cartesian components of the spin operator. It can be shown that the coincidence detection probability provides direct information on $\phi^{(ij)}_{\alpha\beta}(\mathbf{q}_{1},\mathbf{q}_{2};\Omega_{c},\omega_{r})$, which is obtained via Fourier transformation of $\phi^{(ij)}_{\alpha\beta}(\mathbf{q}_{1},\mathbf{q}_{2};t_{c},t_{r})=\phi^{(ij)}_{\alpha\beta}(\mathbf{q}_{1}t_{1},\mathbf{q}_{2}t_{2})$.
Since the two-spin Bethe-Salpeter wave function $\phi^{(ij)}_{\alpha\beta}(\mathbf{q}_{1}t_{1},\mathbf{q}_{2}t_{2})$ describes the time evolution of the magnetic system with two spins excited at times $t_{1}$ and $t_{2}$, the coincidence detection probability effectively probes the dynamical two-spin correlations, resolving both center-of-mass and relative internal dynamics of the spin pairs. As a result, cINS emerges as a promising technique for investigating two-spin correlations in quantum magnetic materials (Su et al., 2021). It holds particular potential for uncovering new physics inaccessible to traditional INS measurements, including the long-sought quantum spin liquids in quantum magnets (Balents, 2010; Zhou et al., 2017; Chamorro et al., 2021).
III.3 cARP/IPES and cARIPES
Coincidence detection of one photoelectric process and one inverse-photoemission process probes two-body correlations of the correlated electrons in the particle-hole channel. — cARP/IPES
Coincidence detection of two inverse-photoemission processes accesses two-body correlations of the correlated electrons in the particle-particle channel. — cARIPES
The cARPES method can be extended to two other coincidence detection techniques, coincidence angle-resolved photoemission and inverse-photoemission spectroscopy (cARP/IPES) and coincidence angle-resolved inverse-photoemission spectroscopy (cARIPES).
In the cARP/IPES measurement, one photon source and one electron source are employed. The incident photon induces a photoelectric process, while the incident electron triggers an inverse-photoemission process, in which it transits into a low-energy state of the target material while simultaneously emitting a photon. The two relevant electron-photon interaction terms governing the photoelectric and inverse-photoemission processes are given by (Su and Zhang, 2020)
$V_{1}$ $=$ $g_{1}d^{{\dagger}}_{\mathbf{k}_{1}+\mathbf{q}_{1}\sigma_{1}}c_{\mathbf{k}_{1}\sigma_{1}}a_{\mathbf{q}_{1}\lambda_{1}},$
$V_{2}$ $=$ $g_{2}c^{{\dagger}}_{\mathbf{k}_{2}\sigma_{2}}a^{{\dagger}}_{\mathbf{q}_{2}\lambda_{2}}d_{\mathbf{k}_{2}+\mathbf{q}_{2}\sigma_{2}}.$ (22)
Here, the creation and annihilation operators are defined as in Eq. (13).
Coincidence detection of the photoelectron emitted in the photoelectric process and the photon emitted in the inverse-photoemission process yields the joint probability of their simultaneous occurrence. The corresponding coincidence detection probability captures the two-body correlations in the particle-hole channel, which can be described by a two-body Bethe-Salpeter wave function defined as (Su and Zhang, 2020)
[…]
More specifically, cARP/IPES provides direct access to its Fourier transformation $\phi^{(3)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1},\mathbf{k}_{2}\sigma_{2};\Omega_{c},\omega_{r}\right)$, where $\Omega_{c}$ and $\omega_{r}$ denote the center-of-mass and inner-pair relative frequencies associated with the particle-hole pairs excited in the relevant photoelectric and inverse-photoemission processes.
cARP/IPES is poised to serve as a powerful tool for elucidating the long-standing puzzle of the microscopic mechanism underlying itinerant magnetism in metallic ferro- and antiferromagnets (Moriya, 1985). In such systems, the magnetic moments originate fundamentally from particle-hole correlations in the spin channel. Moreover, cARP/IPES is expected to contribute significantly to the study of metallic nematic states (Fradkin et al., 2010; Su et al., 2015; Li and Su, 2017), which are governed by particle-hole correlations in the charge channel. It has also been proposed as a promising technique for observing collective Fermi arcs arising from particle-hole excitations in topological Weyl semimetals (Tzeng and Yang, 2020).
In the cARIPES technique, two electron sources are utilized. Electrons from these sources initiate two inverse-photoemission processes, respectively, during which they are injected into the target material, transition to its low-energy electronic states, and simultaneously emit two photons into the vacuum. The electron-photon interaction terms governing these two inverse-photoemission processes are given by
$V_{1}$ $=$ $g_{1}c^{{\dagger}}_{\mathbf{k}_{1}\sigma_{1}}a^{{\dagger}}_{\mathbf{q}_{1}\lambda_{1}}d_{\mathbf{k}_{1}+\mathbf{q}_{1}\sigma_{1}},$
$V_{2}$ $=$ $g_{2}c^{{\dagger}}_{\mathbf{k}_{2}\sigma_{2}}a^{{\dagger}}_{\mathbf{q}_{2}\lambda_{2}}d_{\mathbf{k}_{2}+\mathbf{q}_{2}\sigma_{2}}.$ (24)
Coincidence detection of the two emitted photons reveals the joint probability of the simultaneous occurrence of the two inverse-photoemission processes, which is described by a two-body Bethe-Salpeter wave function in the particle-particle channel (Su and Zhang, 2020)
[…]
Physically, this Bethe-Salpeter wave function captures the dynamical correlated creation of two electrons within the material, thereby characterizing the dynamical correlations of particle-particle pairs. More specifically, the coincidence detection probability in cARIPES directly yields the Fourier-transformed Bethe-Salpeter wave function $\phi^{(4)}_{\alpha\beta}(\mathbf{k}_{1}\sigma_{1},\mathbf{k}_{2}\sigma_{2};\Omega_{c},\omega_{r})$, which distinctly resolves both center-of-mass and inner-pair relative dynamical two-body correlations. Unlike the cARPES case, the excited electrons described by $\phi^{(4)}_{\alpha\beta}(\mathbf{k}_{1}\sigma_{1}t_{1};\mathbf{k}_{2}\sigma_{2}t_{2})$ reside predominantly above the Fermi energy. As a result, cARIPES probes two-body correlations in the particle-particle channel mainly for the electrons above the Fermi energy, making it uniquely suited for investigating Cooper-pair physics in which the paired electrons primarily occupy states above the Fermi energy (Su and Zhang, 2020).
III.4 Coincidence double-tip STS
Coincidence detection of two quantum tunneling currents through two local tips reveals the spatially resolved dynamical two-body correlations of the correlated electrons. — Coincidence double-tip STS
A coincidence detection technique for directly probing spatially resolved dynamical two-body correlations has recently been proposed, based on the double-tip scanning tunneling microscope (STM) and termed coincidence double-tip STS (Su et al., 2025). A schematic illustration of its setup is provided in Fig. 1 of Reference (Su et al., 2025).
The coincidence double-tip STS setup involves two probe tips located at positions $j_{1}$ and $j_{2}$, biased at distinct voltages $V_{1}$ and $V_{2}$ respectively, with each tip carrying a quantum tunneling current. At time $t$, the measured tunneling currents are denoted as $I_{1}(t,n)$ and $I_{2}(t,n)$, where $n$ indexes different measurements. A coincidence tunneling current observable $I_{d}^{(2)}$ is defined as
$I_{d}^{(2)}=\frac{1}{N_{n}}\sum_{n=1}^{N_{n}}I_{1}(t,n)\cdot I_{2}(t,n),$ (26)
with $N_{n}$ being the total number of measurements. Theoretically, $I_{d}^{(2)}$ corresponds to the following statistically averaged correlation function of two local tunneling current operators $I_{1}$ and $I_{2}$ at positions $j_{1}$ and $j_{2}$, respectively, defined by
$I_{d}^{(2)}=\overline{\langle I_{1}(t)I_{2}(t)\rangle}.$ (27)
By introducing a coincidence dynamical conductance
$\sigma_{j_{1}j_{2}}(V_{1},V_{2})=\frac{d}{dV_{2}}\frac{d}{dV_{1}}\overline{\langle I_{1}(t)I_{2}(t)\rangle},$ (28)
a nonequilibrium theoretical framework yields the following expression
$\sigma_{j_{1}j_{2}}(V_{1},V_{2})$ $=$ $-\left(\frac{e^{2}T_{1}T_{2}}{2\pi\hbar^{2}}\right)^{2}\sum_{\sigma_{1}\sigma_{2}}G^{(2)}_{J}(j_{1}\sigma_{1},j_{2}\sigma_{2};\omega_{1},\omega_{2})$ (29)
$\times A_{d_{1}}(0)A_{d_{2}}(0).$
Here, $G^{(2)}_{J}$ denotes a second-order contour-time ordered current correlation function defined by
$G^{(2)}_{J}(j_{1}\sigma_{1},j_{2}\sigma_{2};\omega_{1},\omega_{2})$
$=$ $\iint_{c}dt_{2}dt_{1}\langle T_{c}J_{1A}(j_{1}\sigma_{1};t,t_{1};\omega_{1})J_{2A}(j_{2}\sigma_{2};t,t_{2};\omega_{2})\rangle,$
where $T_{c}$ represents the contour-time ordering operator defined along a closed time contour $t_{i}\rightarrow t\rightarrow t_{i}$, and the integrations are performed along this time contour. The effective current operators $J_{1A}$ and $J_{2A}$ are given by
$J_{1A}(j_{1}\sigma_{1};t,t_{1};\omega_{1})=-ic_{j_{1}\sigma_{1}}(t)c^{\dagger}_{j_{1}\sigma_{1}}(t_{1})e^{-i\omega_{1}(t_{1}-t)}+h.c.,$
$J_{2A}(j_{2}\sigma_{2};t,t_{2};\omega_{2})=-ic_{j_{2}\sigma_{2}}(t)c^{\dagger}_{j_{2}\sigma_{2}}(t_{2})e^{-i\omega_{2}(t_{2}-t)}+h.c..$
The frequencies $\omega_{1}$ and $\omega_{2}$ are related to the bias voltages by $\omega_{1}=-eV_{1}/\hbar$ and $\omega_{2}=-eV_{2}/\hbar$. In Eq. (29), $T_{1}$ and $T_{2}$ are the tunneling constants for the two tunneling currents, while $A_{d_{1}}(0)$ and $A_{d_{2}}(0)$ represent the single-particle spectral functions of the two tip-electron systems at their Fermi energies.
The coincidence dynamical conductance $\sigma_{j_{1}j_{2}}(V_{1},V_{2})$ provides access to dynamical two-body correlations that go beyond the physics captured by traditional single-tip STS technique (Su et al., 2025). For instance, in a nearly free Fermi liquid, the coincidence dynamical conductance reveals two correlated dynamical electron propagation processes: (i) from $j_{1}$ to $j_{2}$ (or vice versa) driven by $V_{1}$, and (ii) from $j_{2}$ to $j_{1}$ (or vice versa) driven by $V_{2}$. In a superconducting state, additional propagation channels originating from the superconducting condensate emerge, coexisting with the normal electron propagation processes.
Experimental implementation of the double-tip STS has advanced considerably over the past two decades (Shiraki et al., 2001; Nakayama et al., 2012; Voigtländer et al., 2018; Leeuwenhoek et al., 2020a, b). However, achieving precise control over tip positionings at extremely short length scales remains a major challenge. Modern double-tip STM setups have successfully realized tip separations on the order of $\sim 30~\text{nm}$ (Nakayama et al., 2012; Leeuwenhoek et al., 2020b). Nevertheless, such resolutions are still insufficient for studying most strongly correlated electron systems, such as various classes of unconventional superconductors (Lee et al., 2006; Chen et al., 2014; Stewart, 2011), where dominant strong correlations are highly localized, with typical correlation lengths comparable to interatomic distances (a few Å). The recent discovery of unconventional superconductivity and Mott physics in magic-angle twisted bilayer graphene (Cao et al., 2018) offers a promising alternative platform for coincidence double-tip STS. In this system, exotic correlated phenomena are inherently connected to the moiré superlattice, which exhibits a characteristic periodicity of $\sim 13~\text{nm}$, a mesoscopic length scale within the resolution of existing double-tip STM setups. Consequently, moiré quantum materials represent an ideal platform for implementing coincidence double-tip STS to probe spatially resolved dynamical two-body correlations in strongly correlated electron systems.
III.5 Double Photoemission
The one-photon-in and two-electron-out scattering processes, with the two emitted electrons coincidentally detected, provide information on the center-of-mass physics of the two-body correlations in the particle-particle channel. — Double photoemission
In 1998, Berakdar formulated the $(\gamma,2e)$ photoemission technique for studying the correlated electrons in solids, now commonly known as double photoemission (Berakdar, 1998). This theoretical framework was subsequently extended to investigate the correlated electrons on solid surfaces (Fominykh et al., 2002) and in various superconducting pairing states (Kouzakov and Berakdar, 2003; Mahmood et al., 2022).
In the double photoemission measurement, a single incident photon excites two correlated electrons from the target material, constituting a one-photon-in and two-electrons-out scattering event. Since the fundamental electrodynamics dictates that a single photon can typically excite only one electron in a photoelectric process, double photoemission must involve additional electron-electron interactions. The relevant interaction terms in the double photoemission measurement comprise the electron-photon interaction $V_{1}$ and the electron-electron interaction $V_{2}$, defined respectively as
$V_{1}(t)=\sum_{\mathbf{k}\mathbf{q}\lambda\sigma}g\,d^{\dagger}_{\mathbf{k}+\mathbf{q}\sigma}(t)c_{\mathbf{k}\sigma}(t)a_{\mathbf{q}\lambda}(t)$ (32)
and
$V_{2}(t_{1},t_{2})$ $=$ $\frac{1}{2}\sum_{\mathbf{k}_{1}\mathbf{k}_{2}\mathbf{q}_{1}\sigma_{1}\sigma_{2}}Uf^{\dagger}_{\mathbf{k}_{1}+\mathbf{q}_{1}\sigma_{1}}(t_{1})e_{\mathbf{k}_{1}\sigma_{1}}(t_{1})$ (33)
$\times d^{\dagger}_{\mathbf{k}_{2}-\mathbf{q}_{1}\sigma_{2}}(t_{2})c_{\mathbf{k}_{2}\sigma_{2}}(t_{2}),$
where the operators $c,d,e,f$ in $V_{2}$ represent the electrons in possibly different energy bands. Here, the interaction constants $g\equiv g(\mathbf{k};\mathbf{q}\lambda)$ and $U\equiv U(\mathbf{k}_{1},\mathbf{k}_{2},\mathbf{q}_{1};t_{1},t_{2})$. It is noted that $U$ generally exhibits strong time-dependent dynamics, which may stem from electron-phonon interactions or spin/charge-density wave fluctuations. For brevity, the band-index dependence of $g$ and $U$ is not explicitly shown.
It is crucial to note that $V_{2}$ does not originate from an externally applied field but arises from the intrinsic electron-electron interactions within the material. Consequently, it can not be directly tuned experimentally. As a result, double photoemission signals are inherently entangled with complex Coulomb drag processes (Devereaux et al., 2023). Although the coincidence detection probability of the double photoemission measurement contains information on the two-body Bethe-Salpeter wave function in the particle-particle channel, it also includes convoluted contributions from summation over inner-pair momenta and/or frequency integration (Mahmood et al., 2022; Su et al., 2024). This entanglement precludes clear resolution of the inner-pair spatial and dynamical physics of the two-body correlations in the target electron system.
A simple explanation for the inability to resolve the inner-pair physics can be given as follows. Suppose the incident photon carries momentum $\mathbf{q}$ and energy $\hbar\omega_{\mathbf{q}}$, while the two emitted photoelectrons have momenta ($\mathbf{k}_{1}^{\prime},\mathbf{k}_{2}^{\prime}$) and energies ($\varepsilon_{\mathbf{k}^{\prime}_{1}},\varepsilon_{\mathbf{k}^{\prime}_{2}}$). Momentum and energy conservations then yield
$\mathbf{q}=\mathbf{k}_{1}^{\prime}+\mathbf{k}_{2}^{\prime}-\Delta\mathbf{k},\,\,\Delta\mathbf{k}=\mathbf{k}_{1}+\mathbf{k}_{2},$
$\hbar\omega_{\mathbf{q}}=\varepsilon_{\mathbf{k}^{\prime}_{1}}+\varepsilon_{\mathbf{k}^{\prime}_{2}}-\Delta E,\,\,\Delta E=E_{1}+E_{2},$ (34)
where ($\mathbf{k}_{1},\mathbf{k}_{2}$) and ($E_{1},E_{2}$) denote the momenta and energies of the two correlated electrons inside the target material. Therefore, in the double photoemission measurement with the fixed parameters $(\mathbf{q},\mathbf{k}_{1}^{\prime},\mathbf{k}_{2}^{\prime},\hbar\omega_{\mathbf{q}},\varepsilon_{\mathbf{k}^{\prime}_{1}},\varepsilon_{\mathbf{k}^{\prime}_{1}})$, the sum of the momenta $\mathbf{k}_{1}$ and $\mathbf{k}_{2}$ and the sum of the energies $E_{1}$ and $E_{2}$ are determined, but the corresponding differences remain unconstrained. This is the fundamental reason why double photoemission can not resolve the internal structure of the two-body correlations in the target electron system.
Nevertheless, since the center-of-mass momentum and energy of the two-body correlations can be accessed via Eq. (34), double photoemission remains a valuable technique for probing the center-of-mass physics of correlated electron pairs. For instance, Mahmood et al. provided a detailed theoretical study of coincidence two-body spectral functions in the pair-density-wave (PDW) and Fulde-Ferrell-Larkin-Ovchinnikov (FFLO) phases, where Cooper pairs possess finite center-of-mass momentum (Mahmood et al., 2022). Double photoemission can also be employed to investigate the collective modes of superconducting condensates (Su et al., 2024), such as the Goldstone-Anderson phase modes (Anderson, 1958), the Higgs amplitude modes (Littlewood and Varma, 1981, 1982), and the Leggett modes in superconductors with two coupled macroscopic superconducting condensates (Leggett, 1966).
IV Discussion and Future Prospects
In the preceding sections, we have briefly explored several recently proposed coincidence detection techniques in different measurement channels. These emerging methods show significant potential for elucidating the breakdown of Wick’s theorem and uncovering interaction-driven many-body correlations in quantum materials (Sobota et al., 2021; Alexandradinata et al., 2025). With the exception of double photoemission, which has been under development for more than two decades, most coincidence detection techniques remain in their infancy. Future advances are anticipated along two primary directions, technological innovation and theoretical development.
On the technological front, a key challenge lies in the limited resolution of current experimental platforms. Addressing this challenge will require continuous, state-of-the-art improvements in both particle sources and detection systems (Sobota et al., 2021; Alexandradinata et al., 2025). For example, enhancing the resolution of coincidence measurements will depend on progress in the controlled emission and/or detection of photons, electrons, or neutrons with high precision in temporal, spatial, energy, and momentum domains. Insights drawn from well-established multiphoton coincidence detection in quantum optics (Strekalov et al., 1995; Pittman et al., 1995; Ferri et al., 2005; Pan et al., 2012) may offer valuable guidance in this context. Moreover, recent rapid advances in attosecond technology (Gaumnitz et al., 2017; Kim et al., 2023; Li et al., 2017; Nob, ) are expected to provide superior photon sources, which will substantially benefit studies of strongly correlated electron systems.
On the theoretical front, a central task is the accurate computation of coincidence detection probabilities, which are directly connected to dynamical two-body correlations in different physical channels. For instance, calculating coincidence detection probabilities for cARPES in the particle-particle pairing channel would provide essential theoretical support for upcoming experiments to study unconventional superconductors. Such calculations could help clarify the microscopic mechanisms of high-Tc superconductivity in cuprate, iron-based, and heavy-fermion superconductors (Anderson, 1987; Lee et al., 2006; Chen et al., 2014; Stewart, 2011, 1984, 2001; Monthoux and Pines, 1993, 1994; Bohrdt et al., 2023; Homeier et al., 2024, 2025). Similarly, theoretical studies of coincidence detection probabilities for cINS would elucidate dynamical two-spin correlations in quantum magnets, potentially revealing new physics linked to quantum spin liquids (Balents, 2010; Zhou et al., 2017). Theoretical predictions for cARP/IPES coincidence detection in the particle-hole channel will be essential for understanding itinerant magnetism in the spin channel (Moriya, 1985) and electronic nematicity in the charge channel (Fradkin et al., 2010). Such efforts may further clarify the roles of antiferromagnetic and multiorbital fluctuations in the emergence of itinerant magnetism and d-wave nematic order in iron-based superconductors (Su et al., 2015; Li and Su, 2017).
It is important to emphasize that the computation of coincidence detection probabilities in different dynamical two-particle channels remains highly challenging, both in theoretical analytical methods and numerical simulations. These challenges originate from the strong many-body correlations in strongly correlated electron systems, where a vast number of degrees of freedom are intrinsically entangled, as well as from the inherent complexity of dynamical two-body response functions. Consequently, new conceptual and methodological advances in theoretical analytical methods and numerical simulations will be essential to unravel the complex and profound physics encoded in coincidence detection probabilities.
Looking ahead, the development of novel coincidence detection techniques specifically tailored for strongly correlated electron systems holds great promise. Moreover, given the strong entanglement among spin, charge, and orbital degrees of freedom in such materials, approaches capable of directly probing correlations between these different degrees of freedom could open entirely new research directions. In particular, the simultaneous detection of two physical processes involving distinct degrees of freedom would allow direct measurement of their mutual correlations and entanglement. Such methods hold the potential to greatly deepen our understanding of strongly correlated quantum matters, especially in shedding light on the complex intertwined quantum phases arising from such strong entanglement.
Acknowledgements
We thank Prof. Jun Chang for invaluable discussions. This work was supported by the National Natural Science Foundation of China (Grant No. 11874318) and the Natural Science Foundation of Shandong Province (Grant No. ZR2023MA015).

### v2

@@
 Systems such as unconventional superconductors (Anderson, 1987; Lee et al., 2006; Chen et al., 2014; Stewart, 2011, 1984, 2001) and quantum magnets (Balents, 2010; Zhou et al., 2017; Chamorro et al., 2021) exhibit a range of unconventional and anomalous phenomena, including non-BCS superconductivity, pseudogap physics, strange metal behavior, electronic nematicity, quantum spin liquids, and quantum phase transitions.
-These emergent properties are governed by intrinsic many-body correlations that can not be adequately described within well established theoretical frameworks like the perturbative Fermi liquid theory or the Ginzburg–Landau-Wilson theory of phase transitions.
+These emergent properties are governed by intrinsic many-body correlations that can not be adequately described within well established theoretical frameworks like the perturbative Fermi liquid theory or the Ginzburg-Landau-Wilson theory of phase transitions.
 The development of new theoretical and experimental methods to study these correlation-driven phenomena remains a significant challenge.
@@
 Second, the existing experimental approaches may be fundamentally limited in their ability to probe the intrinsic many-body correlations and their physical manifestations.
-Unraveling the foundational correlation-driven physics will likely require the development of new experimental theories and advanced technologies.
+Breaking through this impasse and unraveling the foundational correlation-driven physics will likely require the development of new experimental theories and advanced technologies.
+Figure 1: Schematic illustration of coincidence detection.
+(a) Fundamental idea: Coincidence detection of two single-particle perturbation processes reveals dynamical two-body correlations.
+(b) By coincidence detection of two photoelectric processes, coincidence angle-resolved photoemission spectroscopy (cARPES) captures the dynamical two-body correlations in the particle-particle channel.
+(c) By coincidence detection of two neutron-scattering processes, coincidence inelastic neutron scattering (cINS) uncovers dynamical two-spin correlations.
+(d) By coincidence detection of two quantum tunneling currents, coincidence double-tip scanning tunneling spectroscopy (STS) probes the spatially resolved dynamical two-body correlations.
+(e) Coincidence detection technique with post-experiment coincidence counting method.
+Other coincidence detection techniques, such as coincidence angle-resolved photoemission and inverse-photoemission spectroscopy (cARP/IPES), coincidence angle-resolved inverse-photoemission spectroscopy (cARIPES), and double photoemission (DPE), share a similar configuration to (a), but with different incident/scattered particle sources or detectors, for the dynamical two-body correlations in the different particle-particle or particle-hole channel.
 The intrinsic $N$-body correlations in strongly correlated electron systems can be accessed via $N$-body response measurements.
 For example, the intrinsic two-body correlations can be directly probed via two-body response functions which are dominated by the second-order perturbation of the interaction between the target material and an external probe field.
-Guided by this principle, several coincidence detection techniques have recently been proposed for directly measuring two-body correlations in strongly correlated electron systems.
+Guided by this principle, several coincidence detection techniques have recently been proposed for directly measuring two-body correlations in strongly correlated electron systems, as schematically illustrated in Fig.
+1.
 Notable examples include coincidence angle-resolved photoemission spectroscopy (cARPES), which is designed to access two-body correlations in the particle-particle channel, thereby offering a new pathway to elucidate the mechanism of unconventional superconductivity (Su and Zhang, 2020; Devereaux et al., 2023; Cao and Su, 2025).
@@
 III.1 cARPES
-Coincidence detection of two photoelectric processes, in which two incident photons emit two photoelectrons from the target material, captures two-body correlations of the correlated electrons in the particle-particle channel.
-— cARPES
-In cARPES coincidence detection (schematically illustrated in Fig.
-1 (a) of Reference (Cao and Su, 2025)), two incident photons induce two photoelectric processes.
-Two single-photoelectron detectors separately detect the emitted photoelectrons.
+Coincidence detection of two photoelectric processes, in which two incident photons emit two photoelectrons from the target material, captures two-body correlations of the correlated electrons in the particle-particle channel.cARPES
+In cARPES coincidence detection, as schematically illustrated in Fig.
+1 (b), two incident photons induce two photoelectric processes.
 A coincidence detector records the simultaneous occurrence events of two photoelectric processes.
@@
 The coupling constants $g_{1}\equiv g(\mathbf{k}_{1};\mathbf{q}_{1}\lambda_{1})$ and $g_{2}\equiv g(\mathbf{k}_{2};\mathbf{q}_{2}\lambda_{2})$ characterize the interaction strengths.
-The coincidence detection probability for the simultaneous occurrence of two photoelectric processes is given by (Su and Zhang, 2020)
-$\Gamma^{(2)}=\frac{1}{Z}\sum_{\alpha\beta}e^{-\beta E_{\alpha}}\left\lvert\langle\Phi_{\beta}\rvert S^{(2)}(t_{f},t_{i})\rvert\Phi_{\alpha}\rangle\right\rvert^{2},$ (14)
-with initial states $\lvert\Phi_{\alpha}\rangle=\lvert\Psi_{\alpha}\rangle\otimes\lvert 1_{\mathbf{q}_{1}\lambda_{1}}1_{\mathbf{q}_{2}\lambda_{2}}\rangle_{p}\otimes\lvert 0\rangle_{d}$ and final states $\lvert\Phi_{\beta}\rangle=\lvert\Psi_{\beta}\rangle\otimes\lvert 0\rangle_{p}\otimes\lvert 1_{\mathbf{k}_{1}+\mathbf{q}_{1}\sigma_{1}}1_{\mathbf{k}_{2}+\mathbf{q}_{2}\sigma_{2}}\rangle_{d}$.
-Here $\lvert\Psi_{\alpha}\rangle$ and $\lvert\Psi_{\beta}\rangle$ are eigenstates of the target electron system with energies $E_{\alpha}$ and $E_{\beta}$, respectively.
-The relevant second-order $S$-matrix term reads
-$S^{(2)}(t_{f},t_{i})=\left(-\frac{i}{\hbar}\right)^{2}\iint^{t_{f}}_{t_{i}}dt_{2}dt_{1}T_{t}\left[V_{2I}(t_{2})V_{1I}(t_{1})\right].$ (15)
 Let us introduce a two-body Bethe-Salpeter wave function in the particle-particle channel as (Gell-Mann and Low, 1951; Salpeter and Bethe, 1951)
 […]
-In the limit $t_{i}\to$-$\infty$ and $t_{f}\to+\infty$, the coincidence detection probability becomes (Su and Zhang, 2020)
+where $\lvert\Psi_{\alpha}\rangle$ and $\lvert\Psi_{\beta}\rangle$ are eigenstates of the target electron system with energies $E_{\alpha}$ and $E_{\beta}$, respectively.
+In the limit $t_{i}\to$-$\infty$ and $t_{f}\to+\infty$, the coincidence detection probability for the simultaneous occurrence of two photoelectric processes follows (Su and Zhang, 2020)
 $\Gamma^{(2)}=\frac{\left(g_{1}g_{2}\right)^{2}}{\hbar^{4}}\frac{1}{Z}\sum_{\alpha\beta}e^{-\beta E_{\alpha}}\big|\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1},\mathbf{k}_{2}\sigma_{2};\Omega_{c},\omega_{r}\right)\big|^{2}.$ (15)
-Here we have identified $\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1},\mathbf{k}_{2}\sigma_{2};t_{c},t_{r}\right)=\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1}t_{1};\mathbf{k}_{2}\sigma_{2}t_{2}\right)$, where $t_{c}=\frac{1}{2}(t_{1}+t_{2})$ and $t_{r}=t_{2}-t_{1}$ denote the center-of-mass and relative times, respectively.
-The corresponding Fourier transformations are defined by
-$\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1},\mathbf{k}_{2}\sigma_{2};t_{c},t_{r}\right)$
-$=$ $\frac{1}{\left(2\pi\right)^{2}}\iint_{-\infty}^{+\infty}d\Omega d\omega\,\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1},\mathbf{k}_{2}\sigma_{2};\Omega,\omega\right)e^{-i\Omega t_{c}-i\omega t_{r}},$
-$\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1},\mathbf{k}_{2}\sigma_{2};\Omega,\omega\right)$
-$=$ $\iint_{-\infty}^{+\infty}dt_{c}dt_{r}\,\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1},\mathbf{k}_{2}\sigma_{2};t_{c},t_{r}\right)e^{i\Omega t_{c}+i\omega t_{r}}.$
-The center-of-mass frequency $\Omega_{c}$ and the inner-pair relative frequency $\omega_{r}$ in $\Gamma^{(2)}$ are defined as
+Here $\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1},\mathbf{k}_{2}\sigma_{2};\Omega_{c},\omega_{r}\right)$ is the frequency-space Bethe-Salpeter wave function.
+It is the Fourier transformation of $\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1},\mathbf{k}_{2}\sigma_{2};t_{c},t_{r}\right)=\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1}t_{1};\mathbf{k}_{2}\sigma_{2}t_{2}\right)$, where $t_{c}=\frac{1}{2}(t_{1}+t_{2})$ and $t_{r}=t_{2}-t_{1}$ denote the center-of-mass and relative times, respectively.
+The corresponding center-of-mass frequency $\Omega_{c}$ and the inner-pair relative frequency $\omega_{r}$ in $\Gamma^{(2)}$ are defined as
 $\Omega_{c}=\frac{1}{\hbar}(E_{1}^{(2)}+E_{2}^{(2)}),\ \omega_{r}=\frac{1}{2\hbar}(E_{2}^{(2)}-E_{1}^{(2)}),$ (16)
-with $E_{1}^{(2)}$ and $E_{2}^{(2)}$ being the transferred energies in the two photoelectric processes
+where $E_{1}^{(2)}$ and $E_{2}^{(2)}$ are the transferred energies in the two photoelectric processes
 $E^{(2)}_{1}=\varepsilon^{(d)}_{\mathbf{k}_{1}+\mathbf{q}_{1}\sigma_{1}}+\Phi-\hbar\omega_{\mathbf{q}_{1}},$
 $E^{(2)}_{2}=\varepsilon^{(d)}_{\mathbf{k}_{2}+\mathbf{q}_{2}\sigma_{2}}+\Phi-\hbar\omega_{\mathbf{q}_{2}}.$ (17)
-Here $\varepsilon^{(d)}$ are the photoelectron energies, $\hbar\omega$ are the photon energies, and $\Phi$ is the target-material work function.
+Here $\varepsilon^{(d)}_{\mathbf{k}+\mathbf{q}\sigma}$ are the photoelectron energies with momentum $\mathbf{k}+\mathbf{q}$ and spin $\sigma$, $\hbar\omega_{\mathbf{q}}$ are the photon energies with momentum $\mathbf{q}$, and $\Phi$ is the target-material work function.
 It is noted that the Bethe-Salpeter wave function $\phi^{(2)}_{\alpha\beta}\left(\mathbf{k}_{1}\sigma_{1}t_{1};\mathbf{k}_{2}\sigma_{2}t_{2}\right)$ describes the dynamical annihilation of two electrons of the target material.
@@
 More recently, an alternative implementation was proposed (schematically shown in Fig.
-1 (b) of Reference (Cao and Su, 2025)), termed post-experiment cARPES (Cao and Su, 2025).
+1 (e)), termed post-experiment cARPES (Cao and Su, 2025).
 To capture the dynamical two-body correlations in the target electron system, the time interval between the initiations of two photoelectric processes must be shorter than the characteristic time scale $t_{c}$ of the physics under investigation.
@@
 This approach, without the coincidence detector used in the instantaneous cARPES, enables easier and more efficient identification of two-photoelectric-process coincidence events.
-As a result, the post-experiment cARPES technique offers a highly efficient and powerful means for measuring two-body correlations in the target electron system (Cao and Su, 2025).
-It is worth noting that the cARPES technique shares conceptual similarities with time-resolved ARPES measurements of two-electron emission events, which was proposed to investigate electronic noise correlations in materials (Stahl and Eckstein, 2019).
-The related theoretical framework has also been developed to probe two-electron interactions via coincidence detection of two photoelectric processes (Kemper et al., 2025).
+As a result, the post-experiment cARPES technique offers a highly efficient and powerful mean for measuring two-body correlations in the target electron system (Cao and Su, 2025).
+It is worth noting that the cARPES technique shares conceptual similarity with time-resolved correlation ARPES measurements of two-electron emission events (Stahl and Eckstein, 2019).
+Both techniques have the similar fundamental formulations based on perturbation S-matrix theory (Su and Zhang, 2020; Devereaux et al., 2023; Cao and Su, 2025).
+While cARPES measurements probe dynamical two-body correlations in the equilibrium states, time-resolved correlation ARPES measurements investigate these correlations in time-evolution nonequilibrium states.
+By incorporating the temporal envelope of ultrashort photon pulses, time-resolved correlation ARPES measurements enable time-resolved noise two-body correlations in nonequilibrium states.
+This contrasts with cARPES measurements, which provide dynamical two-body correlations in equilibrium states within frequency space, resolving both center-of-mass and relative frequency dependencies.
+More recently, a variant termed correlation ARPES technique has been developed from its time-resolved predecessor (Kemper et al., 2025).
+While it has the same theoretical formulation as the cARPES technique (Su and Zhang, 2020; Devereaux et al., 2023; Cao and Su, 2025), this approach emphasizes the extraction of information on electron-electron interactions in weakly correlated electron systems through the coincidence detection of two photoelectric processes.
 III.2 cINS
@@
 […]
-Following a procedure analogous to that of cARPES, we can introduce the second-order term of the time-evolution $S$-matrix corresponding to the two electron-neutron interaction terms, enabling the calculation of the coincidence detection probability $\Gamma^{(2)}$ for the simultaneous occurrence of two neutron-scattering processes.
 Let us define a two-spin Bethe-Salpeter wave function as (Su et al., 2021; Cao and Su, 2025)
@@
 where $i$ and $j$ label the Cartesian components of the spin operator.
-It can be shown that the coincidence detection probability provides direct information on $\phi^{(ij)}_{\alpha\beta}(\mathbf{q}_{1},\mathbf{q}_{2};\Omega_{c},\omega_{r})$, which is obtained via Fourier transformation of $\phi^{(ij)}_{\alpha\beta}(\mathbf{q}_{1},\mathbf{q}_{2};t_{c},t_{r})=\phi^{(ij)}_{\alpha\beta}(\mathbf{q}_{1}t_{1},\mathbf{q}_{2}t_{2})$.
+It can be shown that the coincidence detection probability $\Gamma^{(2)}$ for the simultaneous occurrence of two neutron-scattering processes provides direct information on $\phi^{(ij)}_{\alpha\beta}(\mathbf{q}_{1},\mathbf{q}_{2};\Omega_{c},\omega_{r})$, which is obtained via Fourier transformation of $\phi^{(ij)}_{\alpha\beta}(\mathbf{q}_{1},\mathbf{q}_{2};t_{c},t_{r})=\phi^{(ij)}_{\alpha\beta}(\mathbf{q}_{1}t_{1},\mathbf{q}_{2}t_{2})$.
 Since the two-spin Bethe-Salpeter wave function $\phi^{(ij)}_{\alpha\beta}(\mathbf{q}_{1}t_{1},\mathbf{q}_{2}t_{2})$ describes the time evolution of the magnetic system with two spins excited at times $t_{1}$ and $t_{2}$, the coincidence detection probability effectively probes the dynamical two-spin correlations, resolving both center-of-mass and relative internal dynamics of the spin pairs.
@@
 $V_{2}$ $=$ $g_{2}c^{{\dagger}}_{\mathbf{k}_{2}\sigma_{2}}a^{{\dagger}}_{\mathbf{q}_{2}\lambda_{2}}d_{\mathbf{k}_{2}+\mathbf{q}_{2}\sigma_{2}}.$ (20)
-Here, the creation and annihilation operators are defined as in Eq. (13).
+Here, the creation and annihilation operators and the coupling constants $g_{1}$ and $g_{2}$ are defined as in Eq. (13).
 Coincidence detection of the photoelectron emitted in the photoelectric process and the photon emitted in the inverse-photoemission process yields the joint probability of their simultaneous occurrence.
@@
 A schematic illustration of its setup is provided in Fig.
-1 of Reference (Su et al., 2025).
+1 (d).
 The coincidence double-tip STS setup involves two probe tips located at positions $j_{1}$ and $j_{2}$, biased at distinct voltages $V_{1}$ and $V_{2}$ respectively, with each tip carrying a quantum tunneling current.
@@
 However, achieving precise control over tip positionings at extremely short length scales remains a major challenge.
-Modern double-tip STM setups have successfully realized tip separations on the order of $\sim 30~\text{nm}$ (Nakayama et al., 2012; Leeuwenhoek et al., 2020b).
-Nevertheless, such resolutions are still insufficient for studying most strongly correlated electron systems, such as various classes of unconventional superconductors (Lee et al., 2006; Chen et al., 2014; Stewart, 2011), where dominant strong correlations are highly localized, with typical correlation lengths comparable to interatomic distances (a few Å).
+Modern double-tip STM setups have successfully realized tip separations on the order of $30~\text{nm}$ (Nakayama et al., 2012; Leeuwenhoek et al., 2020b).
+Nevertheless, such resolutions are still insufficient for studying most strongly correlated electron systems, such as various classes of unconventional superconductors (Lee et al., 2006; Chen et al., 2014; Stewart, 2011), where dominant strong correlations are highly localized, with typical correlation lengths comparable to interatomic distances (a few angstroms).
 The recent discovery of unconventional superconductivity and Mott physics in magic-angle twisted bilayer graphene (Cao et al., 2018) offers a promising alternative platform for coincidence double-tip STS.
-In this system, exotic correlated phenomena are inherently connected to the moiré superlattice, which exhibits a characteristic periodicity of $\sim 13~\text{nm}$, a mesoscopic length scale within the resolution of existing double-tip STM setups.
+In this system, exotic correlated phenomena are inherently connected to the moiré superlattice, which exhibits a characteristic periodicity of approximately $13~\text{nm}$, a mesoscopic length scale within the resolution of existing double-tip STM setups.
 Consequently, moiré quantum materials represent an ideal platform for implementing coincidence double-tip STS to probe spatially resolved dynamical two-body correlations in strongly correlated electron systems.
@@
 Moreover, given the strong entanglement among spin, charge, and orbital degrees of freedom in such materials, approaches capable of directly probing correlations between these different degrees of freedom could open entirely new research directions.
-In particular, the simultaneous detection of two physical processes involving distinct degrees of freedom would allow direct measurement of their mutual correlations and entanglement.
+In particular, the coincidence detection of two physical processes involving distinct degrees of freedom would allow direct measurement of their mutual correlations and entanglement.
 Such methods hold the potential to greatly deepen our understanding of strongly correlated quantum matters, especially in shedding light on the complex intertwined quantum phases arising from such strong entanglement.